why did (7-a)^2 + (2-b)^2 = R^2 become (2-b)^2 - b^2= 0


1
Expert's answer
2021-11-03T11:41:42-0400

"(2-b)^2 + (7-a)^2-R^2=0"


if

"b^2=R^2- (7-a)^2"

then we get

"(2-b)^2 = b^2"